A rectangular plot of land has a length x and a width y, and has a perimeter of 124 m and an area of 945 m2. Find the value of x and y.
Answer by Boreal(15235)

2x+2y=124
x+y=62
xy=945
x=-y+62
substitute
(-y+62)y=945
-y^2+62y=945
0=y^2-62y+945
y=(1/2)+/-(62+/- sqrt (62^2-4(945)); sqrt term is sqrt (64)
y=(1/2)+/- (62+/-8)
y=35, 27 units m
That gives a half perimeter of 62 m and a full perimeter of 124 m
area is 945 m^2